OpenStudy (anonymous):

The seats at a local baseball stadium are arranged so that each row has five more seats than the row in front of it. If there are four seats in the first row, how many total seats are in the first 24 rows? These are the choices 1,476 2,856 2,952 1,428

OpenStudy (anonymous):

מumbers of seats in a row makes an arithmetic series chairs in the nth row an = 4 + 23*5 = 119 and the sum is sn = (4 + 119) * 24/2 = 1476

OpenStudy (anonymous):

ok..I got..thanks for showing work..